Tonya Tonya is 5 3 tall, weighs 151 pounds, and is 38 years old. Her waist circumference is 37. Her life is busy but seden
Answer #1
  • BMI =weight in kg/height in m2

5'3"=160cm

151pounds =151/2.205=68.48

BMI=68.48/2.56=26.75 here over weight range.

<18.5 is under weight

18.5-24.9 healthy weight

25-29.9 is overweight

  • Sedentary woman consuming 1800 calories per day.

BMR IS 68.48*1kcal*24=1651calories

  • 3500calories =1 pounds so cut 500-1000calories a day lose 1-2 pound in a week.
  • Fruits and vegetables such as apple, broccoli, cabbage, asparagus, beet etc
  • soft drinking leads too many calories it can contribute weight gain and obesity. So it must be avoided.